ANNOUNCEMENTS ------------- * HESSI target for today: AR 9866 * PEPCO will perform Preventive Maintenance on a GSFC feeder from March 16 at 5:00 am until March 17 at 6:00 pm. There should be no impact, however as with any outage there is the possibility of power glitches. * There will be a power outage for maintenance work in building 21 on Saturday March 16. FOT REPORT ---------- Spacecraft Status: Nominal Spacecraft Anomalies: 072/2306 BDR 2 on battery 1 switched off unexpectedly. No impacts. Accomplished Activities: VIRGO, SWAN, New Tracking Star (6.5) Planned Activities: VIRGO, NOM HGA Upcoming Operations: None Ground Anomalies: 072/1505 Wayward command. 11 commands retransmitted. 12 min commanding delay. SOLAR STATUS ------------ 03/13 DOY 072 - 03/14 DOY 073 EIT daily On-Line Solar Status and observations: http://umbra.nascom.nasa.gov/eit/plan/log_form20020314.html LASCO daily On-Line Solar Status and observations: http://lasco-www.nrl.navy.mil/observations/2001/Todays_obs.html Web Page for Planning: http://www.bbso.njit.edu/arm INSTRUMENT STATUS ----------------- CDS: Nominal. Today - JOP118 and particle beam study on AR 9866. Friday - JOP118 on AR 9866 and active region mapping study. UVCS: Nominal. Coronal hole observations in the north and mini synoptic today, coronal hole observations and long synoptic Friday. LASCO: Nominal. C2/C3 synoptics. EIT: Nominal. 195 Angstrom CME watch and synoptics. MDI: Nominal. Full disk magnetograms and dopplergrams. After 16 UT MDI will do 4 hours of high res observations as AR 9866 is passing through the high res field of view. TRACE: Nominal. Tracking AR 9866 in 171A and supporting JOP118.